OpenStudy (anonymous):

Simplify the radical √2/3

OpenStudy (austinl):

Is the question, \(\sqrt{\dfrac{2}{3}}\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is that square root over (2/3)

OpenStudy (anonymous):

then multiply and divide by sqrt3 to get a rational number in the denominator

OpenStudy (anonymous):

doing this you will get \[\sqrt{6}/3\]

OpenStudy (austinl):

Ok, \(\sqrt{\dfrac{2}{3}}=\dfrac{\sqrt{2}}{\sqrt{3}}\) You need to multiply both top and bottom by \(\sqrt{3}\). \(\dfrac{\sqrt{2}}{\sqrt{3}}\times \dfrac{\sqrt{3}}{\sqrt{3}}=\dfrac{\sqrt{2}\times \sqrt{3}}{3}=\dfrac{\sqrt{6}}{3}\)

OpenStudy (anonymous):

remember this i used the law \[\sqrt{a*b}=\sqrt{a}*\sqrt{b}\]